### v2

# GCN-DevLSTM: Path Development for Skeleton-Based Action Recognition

Abstract
Skeleton-based action recognition (SAR) in videos is an important but challenging task in computer vision. The recent state-of-the-art (SOTA) models for SAR are primarily based on graph convolutional neural networks (GCNs), which are powerful in extracting the spatial information of skeleton data. However, it is yet clear that such GCN-based models can effectively capture the temporal dynamics of human action sequences. To this end, we propose the G-Dev layer, which exploits the path development – a principled and parsimonious representation for sequential data by leveraging the Lie group structure. By integrating the G-Dev layer, the hybrid G-DevLSTM module enhances the traditional LSTM to reduce the time dimension while retaining high-frequency information. It can be conveniently applied to any temporal graph data, complementing existing advanced GCN-based models. Our empirical studies on the NTU60, NTU120 and Chalearn2013 datasets demonstrate that our proposed GCN-DevLSTM network consistently improves the strong GCN baseline models and achieves SOTA results with superior robustness in SAR tasks. The code is available at https://github.com/DeepIntoStreams/GCN-DevLSTM.
1 Introduction
Action recognition has a wide range of applications in diverse fields such as human-computer interaction, performance assessment in athletics, virtual reality, and video monitoring Liu et al. (2019a). In comparison to video-based action recognition, skeleton-based action recognition (SAR) exhibits better robustness to the variety of light conditions and its efficiency in computation and storage Chen et al. (2021); Shi et al. (2019) while preserving data privacy. Skeleton data can be obtained either by localization of 2D/3D human joints using the depth camera Liu et al. (2019a) or pose estimation algorithms from the videos Cao et al. (2017).
Despite extensive literature devoted to SAR, the challenge still remains open. The key step towards accurate SAR models lies in designing an effective spatio-temporal representation of the skeleton sequence. Previously, Recurrent Neural Networks (RNNs) were commonly used due to their strength in capturing temporal dynamics. The advent of Spatial Temporal Graph Convolutional Networks (STGCN) has shifted the focus towards Graph Convolutional Networks (GCNs), which incorporate both spatial graph convolution and temporal convolution. The majority of subsequent GCN-based approaches Chen et al. (2021); Shi et al. (2020); Cheng et al. (2020b, a); Song et al. (2020); Zhou et al. (2022), following the STGCN framework, focus on the improvement of the spatial graph convolution, with minimal emphasis on improving temporal modules. These methods utilize 1-D temporal convolution operation to capture temporal features. However, whether convolution operation is most effective for capturing temporal dynamics is in question.
Figure 1: The work flow of G-Dev layer.
On the other hand, Rough path theory Lyons (1998) originated as a branch of stochastic analysis to make sense of controlled differential equations driven by highly oscillatory paths. More recently, the applications of rough path theory in sequential data analysis are emerging. In particular, the path signature, as a fundamental object of Rough path theory, has successful applications in various domains ranging from finance to healthcare, demonstrating its capacity as a principled and efficient feature of time series. By embedding time series into the continuous path space, the signature offers a unified treatment to handle variable length and irregular sampling. Signature-based algorithms show promise across diverse fields, including SAR task Liao et al. (2021); Yang et al. (2022); Ahmad et al. (2019); Li et al. (2019a); Cheng et al. (2023). However, the path signature may face challenges such as the curse of dimensionality caused by high path dimension and lack of adaptability due to its deterministic nature Lou et al. (2022). To address these, Lou et al. Lou et al. (2022) propose the path development from rough path theory, as a trainable alternative to path signature. The path development enjoys the desirable properties of the signature while reducing feature dimension significantly. Lou et al. (2022) shows that combining the path development under the appropriate Lie group with LSTM Hochreiter and Schmidhuber (1997) alleviates the gradient vanishing of LSTMs, leading to significant enhancement across various sequential tasks.
This paper places a primary emphasis on enhancing the temporal module within our framework. To this end, we introduce a novel G-Dev layer (as shown in Figure 1) on the temporal graph, which refers to the graph with the node feature as time series. This layer utilizes the sliding window to summarize the local information via the principled path development, offering time dimension benefit and robustness to missing data and irregular sampling. Moreover, we propose the GCN-DevLSTM network, composed of the CNN-based branch and G-Devlopment-based branch for SAR. While retaining the effectiveness of graph convolution for modeling spatial relationships among joints, the GCN-DevLSTM network exploits the hybrid capacity of G-Dev layer and LSTM for handling temporal dynamics. Our GCN-DevLSTM is flexible to incorporate with different GCN backbones. This combination captures discriminative temporal features and improves performance. Experimental results on three commonly used datasets Chalearn 2013, NTU 60 and NTU 120 show that our approach with the GCN backbone being topology-non-shared graph from CTRGCN Chen et al. (2021) outperforms other methods, demonstrating the effectiveness and robustness of our proposed temporal module.
We summarize our main contributions as follows:
  • We design a novel and principled graph development (G-Dev) layer for analyzing temporal graph data by extending the path development originally applied to vector-valued time series. As a generic network on graph data with time-varying node attributes, G-Dev exhibits efficacy and robustness to irregular sampling and missing data.
  • We introduce the novel GCN-DevLSTM network for skeleton-based action recognition that effectively combines the G-Dev layer with LSTM to capture complex temporal dynamics.
  • Experimental results demonstrate that our designed model outperforms other state-of-the-art models in SAR in terms of accuracy and robustness.
2 Related Work
2.1 Skeleton-based Action Recognition
The main challenge in SAR is extracting the effective representation that captures both spatial and temporal dependency of skeleton sequences. Earlier methods typically use convolutional neural networks (CNNs) Chéron et al. (2015); Liu et al. (2017b) and RNN Du et al. (2015); Liu et al. (2017a) to address this challenge without considering structural topology. STGCN Yan et al. (2018) is the first method to capture spatial and temporal relationships using GCN, taking the skeleton topology of the graph into consideration while this topology is not learnable, with only the spatial connection between adjacent joints being connected. As a result, some actions that rely on non-adjacent joints may lead to poor results. Liu et al. (2020) tackles this issue by introducing a learnable adjacency matrix alongside the original manually defined one, enabling connections between non-neighboring joints. Yet, this approach maintains the same graph for all action inputs, which might not be ideal, as different action classes may benefit from different graph topologies. To address this, Li et al. (2019b); Shi et al. (2019, 2020) propose adaptive GCN with the learnable and data-driven topology that shares among all channels. Cheng et al. (2020a); Chen et al. (2021) further propose the non-shared topology for each channel. Current GCN-based methods mainly focus on the enhancement of the graph representation, with limited attention to temporal design. In particular, most GCN-based methods typically use CNNs to capture temporal dynamics. Although studies such as Si et al. (2019); Zhao et al. (2019); Liao et al. (2021) attempt to incorporate LSTM with GCN for the SAR tasks, these methods still significantly underperform the sota GCN models that combine both GCN and CNN on large-scale action datasets.
2.2 Path signature & Path Development
The path signature is a core object of rough path theory, which provides a rigorous way of making sense of the solution to controlled differential equations driven by rough signals Lyons et al. (2007); Lyons (1998). Playing a role as the non-commutative monomials on the path space, the path signature can be used as a faithful feature of time series Lyons (2014); Levin et al. (2013), which leads to applications for various sequential data learning tasks, such as handwritten recognition Xie et al. (2017), writer identification Yang et al. (2015), financial data analysis Lyons et al. (2014), time-series data generation Ni et al. (2021) and SAR Liao et al. (2021); Cheng et al. (2023); Yang et al. (2022); Ahmad et al. (2019); Li et al. (2019a). In the specific context of SAR, Ahmad et al. (2019) focuses on employing path signature solely for extracting spatial features while Yang et al. (2022); Li et al. (2019a) utilize path signature to obtain both temporal and spatial features. However, relying solely on path signature to capture spatial relationships may be suboptimal, as it is not data adaptive and only considers local spatial correlations. Liao et al. (2021) leverages GCN to extract spatial features and employ the log signature to capture temporal dynamics. However, the network tends to overfit rapidly when increasing its depth due to the curse of dimensionality of the signature for the high dimensional path.
More recently, the path development layer Lou et al. (2022) is proposed as a general sequential layer on time series data as a trainable alternative to the signature feature. It has useful applications in supervised learning on sequential data Lou et al. (2022) and synthetic time generation lou2024pcf. This layer is inspired by the (Cantan) development of a path Driver (1995). The path development serves as a useful mathematical tool in rough path theory for proving the properties of the path signature (see e.g.,  Hambly and Lyons (2010); Chevyrev and Lyons (2016)). In particular,Chevyrev and Lyons (2016) proves that under an appropriate Lie group, the path development is a universal and characteristic feature of the path signature, which provides the theoretical support for the development layer Lou et al. (2022). Building upon Lou et al. (2022), our work extends the application of the path development from the vector-valued time series to that of the temporal graph.
3 G-Dev Layer: Development layer of a temporal graph
In this section, we start with a concise overview of the development of the path as the principled feature representation of $d$-dimensional time series. Subsequently, we proceed with introducing the path development layer proposed in Lou et al. (2022). For details of the path development, we refer readers to Lou et al. (2022). Building on this, we propose the G-Dev layer, a novel extension of the development layer applied to temporal graphs, wherein the features of nodes are represented by time series.
3.1 The development of a Path
Let $X\in V([0,T],\mathbb{R}^{d})$, where $V([0,T],\mathbb{R}^{d})$ represents the space of continuous paths of finite length in $\mathbb{R}^{d}$ defined on the time interval $[0,T]$. To fix the notations, let $G$ denote a finite-dimensional Lie group and its associate Lie algebra $\mathfrak{g}$. Examples of the matrix Lie algebra associated with the special orthogonal group (denoted by $\mathfrak{so}$) and special Euclidean group ($\mathfrak{se}$) are given as follows:
$\mathfrak{gl}(m;\mathbb{F})=\{m\times m\text{ matrices over }\mathbb{F}\}\cong\mathbb{F}^{m\times m}\quad(\mathbb{F}=\mathbb{R}\text{ or }\mathbb{C})$
$\mathfrak{so}(m,\mathbb{R}):=\{A\in\mathfrak{gl}(m;\mathbb{R}):A^{T}+A=0\}$
$\mathfrak{se}(m)=\left\{\begin{bmatrix}\omega&v\\ 0&0\end{bmatrix}\mid\omega\in\mathfrak{so}(m,\mathbb{R}),\ v\in\mathbb{R}^{m}\right\}$
Definition 3.1 (Path Development).
Let $M:\mathbb{R}^{d}\rightarrow\mathfrak{g}\subset\mathfrak{gl}(m)$ be a linear map. The path development of $X\in V([0,T],\mathbb{R}^{d})$ on $G$ under $M$ is the solution $Z_{T}\in G$ to the below equation
$dZ_{t}=Z_{t}\cdot M(dX_{t})\quad\text{for all }t\in[0,T]\quad\text{with }Z_{0}=\text{Id}_{m},$ (1)
where $\text{Id}_{m}$ is the identity matrix and $\cdot$ denotes the matrix multiplication.
Example 1 (Linear path).
Let $X\in V([0,T],\mathbb{R}^{d})$ be a linear path. For any linear map $M\in L(\mathbb{R}^{d},\mathfrak{g})$, the development of the path under $M$ admits the explicit formula and is simply
$D_{M}(X)_{0,T}=\exp\left(M(X_{T}-X_{0})\right),$ (2)
where $\exp$ is the matrix exponential.
Thanks to the multiplicative property of the path development (Lemma 2.4 in Lou et al. (2022)), the development of two paths $X\in V([0,s],\mathbb{R}^{d})$ and $Y\in V([s,t],\mathbb{R}^{d})$ under $M$ is the same as the development of the concatenation of these two paths, in formula,
$D_{M}(X*Y)=D_{M}(X)\cdot D_{M}(Y),$ (3)
where $X*Y$ denotes the concatenation of $X$ and $Y$ and $\cdot$ is the matrix multiplication. Eq. (3) enables us to compute the development of any piecewise linear path explicitly.
The path development enjoys several desirable theoretical properties, making it a principled and efficient representation of sequential data. We summarize the following two properties, that are directly relevant to action recognition.
Invariance under Time-reparametrization Similar to the path signature, the path development is invariant under time reparametrization. This property ensures that the development of a path remains unchanged irrespective of the speed at which the path is traversed. It is well suited to action recognition, as the prediction of actions should not depend on the speed of the actions or the frame rate. Time-reparametrization invariance of path development may bring massive dimension reduction and enhance the robustness to variation in speed.
Uniqueness of the path development Intuitively, the path development effectively captures the order of events in time series, leveraging the non-commutative nature of matrix multiplication. Changing the order of events may lead to different development in general, highlighting its capacity to capture the temporal dynamics of sequential data. It is proven in Chevyrev and Lyons (2016) the path development is a characteristic and universal feature of an un-parameterized path in $V_{1}([0,T],\mathbb{R}^{d})$ under the appropriate Lie group. In other words, transforming a path of finite length to its development map $M\mapsto\text{Dev}_{M}(X)$ can uniquely determine the path $X$ up to translation. Thus, the path development combined with the starting point effectively summarizes the path without any loss of information.
3.2 Path Development Layer
Proposed in  Lou et al. (2022), the path development layer is a novel neural network inspired by the path development by exploiting the representation of the matrix Lie group. More specifically, this layer utilizes the parameterized linear map $M\in L(\mathbb{R}^{d},\mathfrak{g})$ by its linear coefficients $\theta\in\mathfrak{g}^{d}$.
Definition 3.2 (Path development layer).
The path development layer, denoted by $D_{\theta}$, is defined as a mapping $\mathbb{R}^{d\times(N+1)}\rightarrow G$, which transforms any input time series $x=(x_{1},x_{2},...,x_{N})$ to the development $z_{N}\in G$ under the trainable linear map $M_{\theta}\in L(\mathbb{R}^{d},\mathfrak{g})$. For $n\in\{0,...,N-1\}$, we set
$z_{n+1}=z_{n}\exp(M_{\theta}(x_{n+1}-x_{n})),z_{0}=Id_{m},$ (4)
where $\theta\in\mathfrak{g}^{d}$ is the model parameter.
Eq. (4) is a direct consequence of Example 1 and the multiplicative property of the development (Eq. (3)). Moreover, Eq. (4) shows the recurrence as an analogy to that of the RNNs, albeit in a much simpler manner, as depicted in Figure 1 (lower panel). Similar to RNNs, the path development can cope with time series input of variable length.
Dimension reduction The output of the path development has dimension $m^{2}$, where $m$ is the matrix size. Note that its dimension does not depend on the path dimension $d$ and time dimension $T$, resulting in the massive dimension reduction. This is in stark contrast to the signature feature of degree $k$, whose dimension $\left(\sum_{i=0}^{k}{d^{i}}=\frac{d^{k+1}-1}{d-1}\right)$ grows geometrically w.r.t $d$. Moreover, the path development is well suited to extract information from high-frequency time series or long time series without concern on the curse of dimensionality.
Robustness to irregular sampling As the path development can be interpreted as a solution to the differential equation (Eq. (1), it is a continuous time series model. Similar to other continuous time models such as signature and Neural CDEs Kidger et al. (2020), the path development exhibits robustness in handling time series data with irregular sampling.
Computational Efficiency of analyzing online data Multiplicative property of the path development allows incremental computation of the development layer in streaming data, thereby reducing both computational load and memory requirements.
3.3 G-Dev Layer: Apply path development layer to a temporal graph
In this subsection, we propose the so-called G-Dev Sequence Layer, which extracts the local information of the temporal graph using the development layer in a rolling window fashion. The core component of the G-Dev Sequence layer serves as a generalization of the development layer from the $d$-dimensional time series to the temporal graph.
Throughout this paper, the temporal graph $\mathfrak{G}_{X}=(\mathcal{V},\mathcal{E},X)$ refers to the graph with each node associated with $d$-dimensional time series features $X$. Here $\mathcal{V}:=\{v^{1},\cdots,v^{M}\}$ denotes the vertex set and $\mathcal{E}$ can be represented by the adjacent matrix $A\in\mathbb{R}^{M\times M}$. We consider the general case where the time dimension of each node may vary across the nodes. More specifically, Let $X^{i}=(x_{t_{1}},\cdots,x_{t_{L_{i}}})\in\mathbb{R}^{d\times L_{i}}$ denote the features of the $i^{th}$ vertex, where $d$ and $N_{i}$ are the path and time dimension, respectively. The feature sets $X=(X^{(1)},\cdots,X^{(M)})$ is obtained by stacking $X^{i}$ for all $i\in\{1,\cdots,M\}$. Clearly, the skeleton sequence is an example of temporal graph data.
G-Dev module. Let us first define the G-Dev module, which is a trainable layer with model parameters $\theta\in\mathfrak{g}^{d}\subset\mathbb{R}^{m^{2}\times d}$, where $m$ is the matrix order of development. It maps a temporal graph $\mathfrak{G}_{X}=(\mathcal{V},\mathcal{E},X)$ to a static graph with matrix-valued features $\mathfrak{G}_{Z}=(\mathcal{V},\mathcal{E},Z)$ by applying the path development to the time-augmented transformation of each feature vector $X_{i}$ of the vertex $v_{i}$, as shown in Figure  1. In formula, for each vertex index $i\in\{1,\cdots,M\}$, the output feature is $Z=(Z^{(i)})_{i=1}^{M}$, where $Z^{(i)}=D_{\theta}(X^{(i)})\in G\subset\mathbb{R}^{m\times m}$.
Built on top of the G-Dev module, we employ a sliding window strategy to construct the following G-Dev Sequence layer, which can summarize (potentially long) time series over coarse time partitions to reduce the time dimension effectively while retaining the fine-grained information by using the low-dimensional representation.
Without loss of generality, let us focus on the case where the graph features $X$ share the same time stamps $(t_{i})_{i=0}^{L-1}$ across vertices [^1] Otherwise, we apply linear interpolation for data imputation, which would not affect the development feature thanks to invariance to time re-parameterization.. Fix the window (kernel) size $K$ and stride $S$. Let $L^{\prime}=\lceil\frac{L-k}{S}\rceil$. By applying padding to $Y$, for any $Y\in\mathbb{R}^{d\times N}$, we define a sequence of time series $\tilde{Y}_{j}$:
$\tilde{Y}_{j}=(Y_{j\times S},\cdots Y_{j\times S+K-1})\in\mathbb{R}^{d\times K},\forall j\in\{0,\cdots,L^{\prime}-1\}.$ (5)
Definition 3.3 (G-Dev Sequence layer).
[…]
The G-Dev Sequence layer (Definition 3.3) can be easily extended to the moving window with the dilation parameter $D$, where we collect the path information at every $D$ time steps during each moving window. To guarantee no information loss of the G-Dev sequence layer, it is crucial to choose $S$ and $K$ to ensure a minimum one-point overlap between two consecutive moving windows.
Figure 2: (a) The pipeline of our proposed approach consists of $N$ blocks, with each block containing a GCN module and a DevLSTM module. (b) The detail of the DevLSTM module.
4 GCN-DevLSTM Network for SAR
[…]
In Figure 2(a), we introduce the proposed GCN-DevLSTM Network, a novel framework for SAR tasks that integrates the GCN and DevLSTM modules into a single block. The GCN module is designed to detect spatial correlations among vertices, while the DevLSTM module handles the temporal dynamics across video frames. Further details on these modules are discussed in Sections 4.1 and 4.2, respectively. Our network is designed to capture discriminative spatial-temporal features with $N$ such blocks, where $N=9$ in this paper. Additionally, within each block, we build a residual path that establishes a shortcut connection with the input to mitigate gradient vanishing/ explosion issues, enhancing the training stability.
Following these GCN-DevLSTM blocks, we employ global average pooling to effectively reduce both temporal and spatial dimensions to 1. After that, a linear function is applied to adjust the feature dimension to match the number of action classes, thereby facilitating the classification process. Cross-entropy loss is employed at the end to compare the predicted action class with the ground truth.
4.1 GCN Module
In the GCN module, we basically follow the network of CTR-GC Chen et al. (2021). However, we depart from the use of three initial adjacency matrices representing self, inward, and outward correlations. Instead, we propose a Hops-CTRGC to enhance the model by incorporating various hops of information to construct three adjacency matrices denoted as $A_{0}$, $A_{1}$, and $A_{2}$.
$A_{0}$ remains an identical (self) matrix. For $A_{1}$, the elements $a_{ij}$ are set to 1 if vertex i and vertex j are physically connected in the skeleton graph; otherwise, $a_{ij}$ in $A_{1}$ is set to 0. $A_{1}$ captures the direct connections in the graph. To enrich spatial information, we introduce $A_{2}=A_{1}^{2}$, enabling a vertex to connect with its neighbour’s neighbour. $A_{2}$ allows us to extend our understanding beyond direct connections to include relationships involving two hops in the graph. Since GCN module is not our main focus in this paper, readers can find more details about the network of CTR-GC in Chen et al. (2021) or in our supplementary material 7.5.
4.2 G-DevLSTM Module
The proposed DevLSTM module, illustrated in Figure 2(b) is mainly composed of the CNN-based branch and the G-development-based branch. The output of DevLSTM module is given by applying GeLU activation function to the summation of the output of both branches.
G-Development branch. Inspired by the design of multi-scale TCN Liu et al. (2020), we also design a multi-scale network to capture temporal dynamics. We utilize two G-dev layers with the dilation parameter $D=1$ and $D=2$ to enhance its ability to capture temporal patterns across different scales in the input sequence.
The path development combined with the starting point can determine the path uniquely. Hence, we append the location of the starting points to the output of the path development layer along the feature dimension over each sub-time interval. The resulting features are then fed into an LSTM network, which is able to effectively aggregate the local development summary and learn deep features to extract the long-term dependencies in time series data. We use the sequential output of LSTM, which keeps the time dimension the same as the input.
CNN-based branch. MaxPool and a convolutional layer with $5\times 1$ kernel size have consistently demonstrated superior performance within the temporal module of other GCN-based studies. Introducing both MaxPool and the $5\times 1$ convolutional operation to our DevLSTM temporal module has the potential to enhance model performance even further. We can reasonably design kernel size, padding, stride and dilation to ensure the output size of Maxpool, LSTM and $5\times 1$ convolution remain the same.
5 Numerical Experiments
Table 1: Experimental results on NTU-60 and NTU-120 dataset. The best results for each benchmark are bolded.
Venue NTU-60 NTU-120
X-sub (%) X-view (%) X-sub (%) X-view (%)
ST-GCN Yan et al. (2018) AAAI’18 81.5 88.3 70.7 73.2
AGC-LSTM Si et al. (2019) ICCV’19 89.2 95.0 - -
Bayesian-LSTM Zhao et al. (2019) ICCV’19 81.8 89.0 - -
2S-AGCN  Shi et al. (2019) CVPR’19 88.5 95.1 82.5 84.2
MS-AAGCN  Shi et al. (2020) TIP’20 90.0 96.2 - -
Shift-GCN  Cheng et al. (2020b) CVPR’20 90.7 96.5 85.9 87.6
Logsig-RNN Liao et al. (2021) BMVC’21 - - 75.8 78.0
CTR-GCN(J) Chen et al. (2021) ICCV’21 89.8 94.8 84.9 86.7
CTR-GCN(J+B) Chen et al. (2021) ICCV’21 92.0 96.3 88.7 90.1
CTR-GCN Chen et al. (2021) ICCV’21 92.4 96.8 88.9 90.6
EfficientGCN Song et al. (2022) TPAMI’22 92.1 96.1 88.7 88.9
Ta-CNN Xu et al. (2022) AAAI’22 90.4 94.8 85.4 86.8
FR-Head Zhou et al. (2023) CVPR’23 92.8 96.8 89.5 90.9
Our methods Joint Bone Dual Joint motion Bone motion
90.8 95.7 86.2 87.7
91.1 95.2 87.0 88.6
91.2 95.3 87.1 88.7
92.4 96.4 89.0 90.4
92.6 96.6 89.3 90.7
92.9 96.6 89.7 91.2
To validate the effectiveness of the proposed approach, we carry out numerical experiments on the following popular benchmarking datasets for skeleton-based action and gesture recognition. These datasets are chosen for their diversity in data size and task complexity, providing a comprehensive evaluation of the model’s performance. (1) Chalearn2013 Escalera et al. (2013) is a public dataset for gesture recognition, consisting of 11,116 skeleton examples and 20 gesture categories, filmed by 27 people. Each frame has 20 joints. (2) NTU RGB+D 60 (NTU-60) Shahroudy et al. (2016) consists of 56,880 action samples and 60 action classes, filmed by 40 people using three Kinect cameras in different camera views. 3D skeletal data in the dataset have different sequence lengths, with each owning 25 joints. (3) NTU RGB+D 120 (NTU-120) Liu et al. (2019b) extends upon NTU RGB+D 60, consisting of 114,480 samples and 120 action classes, filmed by 106 people in three camera views. We consider the Cross-subject (X-Sub) and Cross-view (X-View) evaluation benchmarks for both NTU-60 and NTU-120 as proposed by Shahroudy et al. (2016). We use only skeleton data from these datasets for our experiments.
Experiments are conducted using a single Nvidia A6000 graphics card, where a batch of 16 examples is fed to the network. Stochastic gradient descent optimiser Bottou (2012) is employed, with a Nesterov $\text{momentum}=0.9$ and a weight decay $=0.0003$. The training epoch is set to 65 and we apply the warm-up strategy He et al. (2016) to the first 5 epochs. The initial learning rate is 0.02 and a step learning rate decay is applied with a factor of 0.1 at epoch 35 and 55, respectively. Data processing on NTU datasets mainly follows the procedure in Chen et al. (2021), with the length of all input data sequences resizing to 64. We use $\mathfrak{se}$ Lie algebra in this paper, with a corresponding matrix size of $m=10$. A detailed network architecture can be found in the supplementary. The replicate padding is used in our experiments.
5.1 Fusion results from different data streams
Most of the recent approaches that we are set to compare combine predicted scores from different input data streams for enhanced performance. The most commonly used data streams are joint (J), bone (B), joint motion (JM) and bone motion (BM) where details can be found in Shi et al. (2019, 2020). Apart from the above-mentioned data streams, we introduce a novel stream: the dual graph (D). In a dual graph, vertices represent bones, and edges establish connections between these bones. More details about dual graph can be found at Supplementary 7.1.
5.2 Comparison with state-of-the-art methods
Table 2: Results on Chalearn2013 dataset.
Acc (%)
ST-LSTM + Trust Gate Liu et al. (2018) 92.00
Multi-path CNN Liao et al. (2019) 93.13
3s_net_TTM Li et al. (2019a) 92.08
GCN-Logsig-RNN Liao et al. (2021) 92.86
CTR-GCN Chen et al. (2021) 95.19
ST-PSM+L-PSM Cheng et al. (2023) 94.18
Our method 95.61
We compare our approach with other state-of-the-art methods on both NTU-60 and NTU-120 datasets, as shown in Table 1. Since our GCN module is following CTR-GCN Chen et al. (2021) with the major difference lying in the temporal module, we show more results on CTR-GCN for a more comprehensive comparison. As we can see all of our results, no matter whether using only joint or using ensembled streams combining joint and bone, outperform CTR-GCN with a large margin on all benchmarks. Our method’s limited performance gain with extra joint or bone motion, compared to CTR-GCN, may stem from our path development layer exploiting position differences between consecutive time points. As a result, streams using joint and bone motion provide identical information at the beginning of the network, limiting their help. Our model, integrating five data streams, achieves the best result compared to other state-of-the-art approaches across three out of four benchmarks.
Figure 3: Robustness analysis on NTU60 X-sub benchmark.
Table 2 shows the comparison results between our result and other state-of-the-art methods on Chalearn2013 dataset. We only employ the joint stream for fair comparison as other approaches only use the joint representation on Chalearn2013 dataset. It is evident that our model yields the best result on the small-scale gesture dataset, proving its effectiveness on the small dataset. It is important to highlight that our network structure remains the same on all datasets, demonstrating its strong adaptability for different dataset sizes.
5.3 Comparison with signature-based models
Compared to the Logsig-RNN Liao et al. (2021), one of the recent signature-based action recognition methods (using the joint stream only), our model (J) outperforms 10.4% and 9.7% respectively on both benchmarks of NTU-120 dataset, and 2.75% on Chalearn2013 dataset. Compared to the latest signature-based action recognition approach Cheng et al. (2023), we still outperform 1.43% on Chalearn2013 dataset while unfortunately, they didn’t report their results on NTU dataset, leaving no way to compare. These results support the advantages of using path development that we claimed before, compared with path signature.
5.4 Comparison with MS-TCN
To demonstrate the flexible integration of our DevLSTM temporal module with various GCN modules in a plug-and-play fashion, we employ the adaptive graph convolution (AGC) graph in 2S-AGCN Shi et al. (2019) and a fixed graph without learnable parameters, respectively as an additional GCN module backbone in our network. In addition, we compare our DevLSTM with multi-scale temporal convolution network (MS-TCN) used in Chen et al. (2021) across the different GCN backbones. Table 3 illustrates that the proposed DevLSTM module consistently outperforms MS-TCN with different GCN backbones by a large margin, indicating that our designed temporal module is effective in enhancing various GCN backbones. An overview of adaptive graph convolution and fixed graph can be found in the supplementary material.
Table 3: Comparison between our DevLSTM module with MS-TCN across various GCN modules on NTU60 dataset.
CTR-GC MS-TCN (J) 90.3 94.9
DevLSTM (J) 90.8 95.7
AGC MS-TCN (J) 89.8 94.5
DevLSTM (J) 90.6 95.6
Fixed graph MS-TCN (J) 88.7 93.1
DevLSTM (J) 89.5 94.5
Table 4: Ablation studies of DevLSTM module on NTU-60 dataset. w/o means without.
Components X-sub X-view
with all components (J) 90.8 95.7
w/o residual path (J) 89.9 95.2
w/o 1$\times$ 1 conv & 5$\times$1 conv (J) 90.7 95.3
w/o 1$\times$ 1 conv & 3$\times$ 1 maxpool (J) 90.6 95.4
w/o path dev (J) 90.5 95.2
5.5 Ablation studies
We further investigate the importance of residual path design to connect the input and output of each block and each component in our temporal module. As evident from Table 4, residual path plays an important role to enhance the model performance since it can help alleviate gradient vanishing/explosion in the deep neural network. For temporal components in our DevLSTM module, the path development layer emerges as the primary contributor to the overall performance improvement since removing it causes the largest performance degradation, with other components also contributing to enhanced performance.
5.6 Robustness Analysis
We comprehensively compare the robustness of our full model (GCN-DevLSTM), CTRGCN Chen et al. (2021) and our model without path development layer (GCN-LSTM) in terms of missing frames, missing joints and variable sequence length. To this end, we firstly create a new test dataset by randomly dropping or inserting a certain number of frames out of 64 frames to test the robustness of missing frames and variable length. Figure 3a illustrates that our full model with the path development layer consistently outperforms the other methods in terms of accuracy, regardless of the proportion of frames missing or inserting, proving its robustness to irregular sampling and variable sequence length.
Additionally, to validate the robustness of missing joints due to occlusions in real life, we create another dataset by randomly dropping a certain number of joints in each frame. To fill in these missing joints, two techniques are adopted: zero-filling (in Figure 3b) and linear interpolation (in Figure 3c). Both figures show that our model with G-Dev layer, is consistently more robust against missing joints than others. More details are in the supplementary material.
6 Conclusion and future work
Conclusion. In summary, we propose a novel and general GCN-DevLSTM network for analyzing temporal graph data by leveraging the path development layer. This approach not only offers flexibility when combined with various GCN modules but also significantly enhances performance by effectively capturing the temporal dependencies within data. The effectiveness of the proposed method is validated by the state-of-the-art results on both small gesture data and large action datasets in SAR. Moreover, it demonstrates superior robustness to missing data and variable frame rates. Limitation. The current G-Dev Sequence layer relies on other GCNs to extract spatial information, which limits its standalone utility. Future enhancements could include integrating the features of adjacent nodes directly into the input of the path development layer to better capture spatio-temporal dependencies. Broader impact Our proposed G-Dev Sequence layer provides a generic model for temporal graph data, opening up possibilities for its applications across various fields, including urban planning, social network analysis and others.
Acknowledgements. LJ, WY and HN are supported by the EPSRC [grant number EP/S026347/1]. WY and HN are also supported by The Alan Turing Institute under the EPSRC grant EP/N510129/1.